If the average (arithematic mean) of 5 consectve integers is 12,what is the sum of the least and greatest of the 5 integers?
Answer by Boreal(15235)

average=total/n
n*average=total
5*average=total
lowest number is x
greatest number is x+4
One can add the numbers from x to x+5. One may also multiply the middle number, x+2, by 5. The middle number is 12, and 12*5=60.
x+x+1+x+2+x+3+x+4=60

5x+10=60
x=10 SMALLEST
x+4=14 LARGEST THEIR SUM (WHICH IS WHAT IS DESIRED) IS 24.
10+11+12+13+14=60.